OpenStudy (anonymous):

write the slope for m= 3/2;(7,4)

OpenStudy (anonymous):

??? the slope is given do you mean write the equation of the line?

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

take your point and slope, substitute them into the point slope equation, and solve for either standard form or slope intercept

OpenStudy (anonymous):

i dont know how to do stuff like that yet

OpenStudy (anonymous):

look at the point slope formula see x1 and y1? (x1,y1)= (7,4) x1=7 y1=4 m=3/2

OpenStudy (anonymous):

ok yea i see

OpenStudy (anonymous):

so i divide 3 and 2 then what?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

no leave the slope as a fraction

OpenStudy (anonymous):

you know x1, y1, and m so "plug them into" the point slope formula above

OpenStudy (anonymous):

so x1 means 7 and y1 means 4 so my answer is m=3/2 ?

OpenStudy (anonymous):

no m is the slope....

OpenStudy (anonymous):

oh lol so what do i do to get my answer

OpenStudy (anonymous):

we want the equation of the line, not the slope slope is already given....

OpenStudy (anonymous):

read through my crap, until you think you fully understand it then send me a message

OpenStudy (anonymous):

sorry im trying to understand it

OpenStudy (anonymous):

y-y1=m(x-x1) y1=7 m=3/2 x1=4 we substitute them into the point slope equation... y-7=3/2(x-4) how do you not understand substitute

OpenStudy (anonymous):

i didnt have any of that when i went to school lol

OpenStudy (anonymous):

... grade ?

OpenStudy (anonymous):

school since 1993 then got mu high school diploma at home , now im trying to do college and i pass everything except algebra

OpenStudy (anonymous):

you shouldve been taught algebra in middle school if a=b and b=c then a=c yes or no

OpenStudy (anonymous):

no clue sorry thats why im trying to read everything but its hard without a teacher lol...

OpenStudy (anonymous):

if

OpenStudy (anonymous):

if A is equal to B B is equal to C is A equal to C?

OpenStudy (anonymous):

lol my son said he didnt know neither

OpenStudy (anonymous):

this is true it is known as the transitive property and one of the basis of substitution

OpenStudy (anonymous):

i will remember that lol thanks

OpenStudy (anonymous):

do you know what a variable is?

OpenStudy (anonymous):

changing something/

OpenStudy (anonymous):

not exactly in algebra/ math a variable is a symbol (usually letters) that represent a number but is not actually one think of it kind of like an empty chair, it represents that a person is sitting there but the person is not actually there sitting in it

OpenStudy (anonymous):

oh kool

OpenStudy (anonymous):

commonly used variables are x, y,a,b,c now where im saying x=2 im giving the x variable a value in this case, x is equal to two (filling up that chair with 2 pieces of paper) if i say x=7y it means that x is equal to 7 times another variable analogy an empty chair holds the same amount as 7 empty buckets

OpenStudy (anonymous):

now example if i gave you this expression x+y=z and also said x=1 and y =2 and wanted you to find z do you think you can do it?

OpenStudy (anonymous):

1+2=3

OpenStudy (anonymous):

basically what we do here is we substitute we dont know the value of z but we know the value of x and y therefore we will get the equation 1+2=z z=3

OpenStudy (anonymous):

yay lol

OpenStudy (anonymous):

ok now what if i said x=3 and y=x-1 find z for x+y=z?

OpenStudy (anonymous):

since we know y is equal to x-1, we can substitute that in place of y so we will then get x+(x-1)=z because we alreayd know x is equal to 3 3+(3+1)=z